# Spatially resolved broad line region in a quasar at z=4

### v1

@@
We present the first near-infrared interferometric data of a QSO at z=4. The K-band observations were performed with GRAVITY+ on the VLTI using all 4 UTs, detecting a differential phase signal that traces the spatially resolved kinematics for both the H and H lines in the broad line region. We fit the two lines simultaneously with an updated model that includes distinct rotating and conical outflowing components. We find that more than 80% of the H I line emission from the BLR originates in an outflow with a velocity up to km s-1. This is oriented so that our line of sight is along an edge of the conical structure, which produces the prominent blue wing on the line profile. A combination of anisotropic line emission and mid-plane opacity lead to the single-sided phase signal. The model is able to qualitatively match both the outflowing C IV line profile and the systemic O I fluorescent emission. The derived black hole mass of M⊙ is the highest redshift black hole mass measurement to date obtained directly from BLR dynamics. It is an order of magnitude lower than that inferred from various single epoch scaling relations, and implies that the accretion is highly super-Eddington. With reference to recent simulations, the data suggest that this QSO is emitting close to its radiative limit in a regime where strong outflows are expected around a polar conical region.
@@
Galaxies: active – quasars: individual: J0529 – quasars: emission lines – quasars: supermassive black holes – Infrared: galaxies1 Introduction
@@
The cosmic epoch at marks a key period between the end of reionisation at (Robertson, 2022; Loeb & Barkana, 2001) and the most active phase of galaxy formation at cosmic noon around (Förster Schreiber & Wuyts, 2020; Tacconi et al., 2020). There is an increasing interest in spatially resolving the molecular and ionised emission lines in galaxies on kiloparsec scales during this epoch (Le Fèvre et al., 2020; Wylezalek et al., 2022; Galbiati et al., 2023; Herrera-Camus et al., 2025; Übler et al., 2024), because doing so can provide insights into the processes that have shaped the universe into what it is now. The role of active galactic nuclei (AGN) is equally important because the black hole accretion rate is rising rapidly during this epoch until it peaks around the same time as galaxy formation (Aird et al., 2015; Kim et al., 2024). Indeed, both the major impact that feedback from AGN can have on galaxy growth (Fabian, 2012; King & Pounds, 2015; Fan et al., 2023), and the existence of local relations between black hole mass and their host galaxy properties (Ferrarese & Merritt, 2000; Gebhardt et al., 2000; Gültekin et al., 2009; McConnell & Ma, 2013; Saglia et al., 2016), indicate that supermassive black holes should co-evolve with galaxies (Kormendy & Ho, 2013; Heckman & Best, 2014).
@@
Observations with JWST, as nicely illustrated by Adamo et al. (2025), have recently highlighted a population of AGN with luminosities overlapping those of local Seyfert nuclei, spanning the entire redshift range from cosmic noon to reionisation (Kocevski et al., 2023; Übler et al., 2023; Greene et al., 2024; Maiolino et al., 2024; Matthee et al., 2024; Suh et al., 2025; Juodžbalis et al., 2025; Scholtz et al., 2025). The data have unexpectedly indicated that, compared to local relations, these black holes are overmassive with respect to their host galaxies. The implication is that they must have grown very rapidly, requiring sustained near-Eddington, or even super-Eddington, accretion, depending on the mass of the seed black holes (Kokorev et al., 2023; Larson et al., 2023; Bogdán et al., 2024; Jeon et al., 2024; Maiolino et al., 2024; Mehta et al., 2024; Suh et al., 2025). However, concerns have been raised about the scale of uncertainties on both the black hole mass and host galaxy mass estimates, and on the use of scaling relations to derive black hole masses (Lambrides et al., 2024; Lupi et al., 2024; Adamo et al., 2025; Bertemes et al., 2025).
@@
Complementary to these efforts are studies focussing on highly luminous QSOs, which are expected to drive the powerful outflows that have the greatest impact on host galaxy evolution (Bischetti et al., 2017). Such QSOs are inferred to have black hole masses M⊙, with a significant fraction close to M⊙, and so are radiating at around their Eddington luminosity (Vietri et al., 2018). The mass outflow rates derived from the prominent broad [O III] lines exceed 1000 M⊙ yr-1, and have kinetic energies around 1% of the bolometric luminosity. Vietri et al. (2018) showed that they follow the expected trend (Sulentic et al., 2011; Sulentic & Marziani, 2015), where those exhibiting prominent [O III] have moderately strong C IV emission with modest blue-shift, while QSOs without such strong [O III] emission have weaker C IV that is more strongly blue-shifted. These authors argued that this could be due to orientation effects affecting observation of the outflows in the broad line region (BLR); and comparison of the C IV blue-shift and FWHM led them to propose that the line profile is a combination of virialised and polar outflowing components. Extremely luminous QSOs at have also been targeted by mid-infrared JWST observations with the aim of resolving the host galaxy (Wylezalek et al., 2022; Vayner et al., 2024), The data have revealed extended [O III] emission in the host galaxies, the outflows, and the circumgalactic medium.
@@
While the properties of outflows on large scales correlate with AGN luminosity, there are also significant departures, for example due to the short timescale variability of AGN which can lead to a wide range of mass loading factors (Zubovas & Nardini, 2020). Similarly, how the outflow is driven, the relative importance of different gas phases, the role of dust, and the degree of clumpiness, all have an impact on the outflow properties and may change with distance from the AGN (King et al., 2011; Faucher-Giguère & Quataert, 2012; Zubovas & King, 2012; Somerville & Davé, 2015; Ward et al., 2024). As such, resolving structures in these massive outflows nearer to their launching sites is a central part of understanding them, but doing so requires additional techniques due to the extremely small scales involved. Velocity resolved reverberation mapping (RM) of nearby AGN has shown that there can be an outflow component to the BLR (Denney et al., 2009; Grier et al., 2013; Du et al., 2016a). More recently, near-infrared interferometry with GRAVITY has been used to spatially resolve the BLR (GRAVITY Collaboration et al., 2018, 2020, 2021), finding that in some local AGN, outflows can dominate the BLR kinematics (GRAVITY Collaboration et al., 2024). The first QSO at has now been observed with the same technique, resolving the BLR kinematics in a highly accreting QSO (Abuter et al., 2024). Observations of several more luminous QSOs at the same epoch have revealed clear signatures of radial flows in a significant fraction of their BLRs (GRAVITY+ Collaboration et al., in prep).
@@
In this paper we take the next step towards higher redshifts, and present interferometric data from GRAVITY+ that resolve the BLR in SMSS J052915.80-435152.0 (hereafter J0529), a highly luminous QSO at z=3.96 (Wolf et al., 2024). Our focus is two-fold. One aspect is to probe the AGN driven outflow on sub-parsec scales, where a broad-line region outflow may be launched. The other is to make a dynamical estimate of the black hole mass, as part of a larger effort to understand whether the scaling relations derived locally can be reliably applied at high redshifts and if needed re-calibrate those relations. Our observations are presented in Sec. 2. We provide a context for our analysis in Sec. 3 by reporting the BLR size and BH mass expected from scaling relations. We then describe our BLR model in Sec. 4 and present our fitting results in Sec. 5, speculating about why they differ so significantly from the scaling relations, and showing how the model can be used to match other emission lines. We summarise our results in Sec. 6.
@@ 2 Observations and Data Reduction
The data analysed here were obtained with GRAVITY (GRAVITY Collaboration et al., 2017) as part of programme 114.27UV. They make use of the dual field mode of GRAVITY-wide (GRAVITY+ Collaboration et al., 2022) and the new adaptive optics modules (GRAVITY+ Collaboration, 2025), which are upgrades towards GRAVITY+. As used here, GRAVITY combines the light of all four 8-m UTs of the VLT, to probe spatial scales of milli-arcsec and perform spectro-astrometry at 10 micro-arcsec levels.
The data were obtained on five nights from Sep. to Dec. 2024 with typical seeing (at zenith, and at 500 nm) on different nights of 0.5″ and 0.7″, and coherence timescales of ms. The MEDIUM-COMBINED dual-field wide mode was used, meaning that medium spectral resolution () was used, the polarisations were combined, and the science target and fringe tracking reference star were separate sources with more than 2″ between them. A total of almost 7.8 hr integration was obtained in 64 exposures as shown in Table 1. After processing the data with the GRAVITY instrument pipeline (v. 1.7.0b1), particularly noisy data sets were rejected, reducing the number used from 64 to 59. An additional step was also made to compensate the wavelength dependent variability caused by the dispersion of air in the non-vacuum delay lines as described in Appendix A of GRAVITY Collaboration et al. (2020), using updated components calibrated from GRAVITY-wide observations. Due to the limited signal-to-noise of individual science frames, it was applied on a night-by-night basis: the scaling derived from the combined frames on each night (excluding spectral regions near the H and H lines) was applied to each of those frames individually. Although this approximation cannot take into account changes in delay line length due to tracking, it appears to be sufficient in these circumstances. The resulting differential phase spectra from all nights were then combined with equal weighting, after rejecting (at most a few) deviant values in each spectral channel.
@@
27 Sep 2024 | 0.47 | 8.4 | 3 | 4300 |
14 Nov 2024 | 0.72 | 5.9 | 20 | 4100 |
15 Nov 2024 | 0.48 | 10.4 | 22 | 4100 |
12 Dec 2024 | 0.53 | 5.9 | 6 | 4100 |
14 Dec 2024 | 0.72 | 4.7 | 13 | 4100 |
To support the GRAVITY data, an additional K-band spectrum was obtained with the ERIS adaptive optics imager and spectrometer (Davies et al., 2023) as part of programme 112.25M3. ERIS includes an integral field spectrometer (SPIFFIER) which was used here in seeing limited conditions (0.71″ at zenith and 500 nm) with the largest 8″ field of view. The low resolution K-band grating covers 1.96-2.45 m at a resolution of . The data were taken as 11 dithered exposures of 300 s, yielding a total integration time of 55 mins, and reduced with the ERIS instrument pipeline (v 1.6.4). Telluric correction and flux calibration were performed using the B9 IV standard star HD 38343 which has K=7.87 mag. Following this, additional steps were performed to interpolate over the noisiest pixels in the datacube and reduce wavelength dependent variations in the background level. The flux density of the QSO was measured from a spectrum extracted in a large 1.25″ aperture, yielding K = 13.88 mag.
@@ 2.1 Image
An image was extracted from the datacube by summing channels in the range 2.028–2.290 m, indicating that the spatial resolution achieved on the QSO itself in the K-band was 0.43″. The image is shown in Fig. 1 on a logarithmic scaling to highlight faint sources. The QSO shows no sub-structure, and the only other source in the field is the M dwarf star previously noted by Wolf et al. (2024) a few arcsec north of the QSO. This strengthens the assessment made by those authors that there is no sign of strong lensing, and hence the luminosity is intrinsic.
@@ 2.2 Spectral decomposition
A final spectrum with higher signal-to-noise (S/N) of 85 (but slightly lower flux) than that above was extracted in a 0.62″ aperture better matched to the FWHM. This is used to recover the spectral profile of the broad H I lines, which requires decomposition of the spectrum in order to deblend them from the Fe II features which are known to be strong in highly accreting QSOs (Boroson & Green, 1992; Du et al., 2016b). Since it is already clear from Fig. 2 that the H I lines have a prominent blue wing, we fit these with a pair of Gaussians. This is done solely to match the overall profile and we do not consider the Gaussians individually. We use the same velocity profile for both H I lines, which, although not physically required, is necessary due to limitations of the data. In addition we required zero velocity offset between the H I lines, which was an important constraint on the fit. The reason is that the blue wing of the H line is strongly blended with Fe II and so is primarily constrained by the H line. On the other hand the red side of the H line does provide a primary constraint, whereas the H line is affected by both atmospheric absorption (leading to narrow features and also affecting the slope) and potentially some emission from a blue-shifted wing of [O III] line emission, the peak of which would be just out the spectral range. While we do not have direct evidence for [O III] (and indeed the [O III] line at 4363 Å is not detected although it is expected to be a factor 10-100 weaker), in luminous high-z samples it can be associated with an outflow as indicated by the strong blue-shifted C IV emission described in Sec. 3.2 (Marziani et al., 2017). The intrinsic continuum is expected to be a power law, but over this narrow wavelength range we have adopted a linear function. We consider several templates for the optical Fe II emission, including those based on I Zw 1 (Véron-Cetty et al., 2004), Mrk 493 (Park et al., 2022), a composite based on several QSO spectra (Tsuzuki et al., 2006), and an adaptable template derived from multiple AGN (Kovačević et al., 2010). After trying these options, we adopted the template from Park et al. (2022) because it provides the best match to the Fe II emission on the blue side of the H line; in other regions the templates were equivalently good. Although the properties of the Fe II complex were not tied to the Balmer lines, the broadening of km s-1 FWHM is comparable to the width of the Balmer lines as one would expect if they arise in similar locations (Hu et al., 2015). Finally, specific features in the spectrum that were clearly due to atmospheric effects were excluded during the fitting process. The resulting fit has , and would imply a modest update to the redshift of J0529 as based on the line peak – but see Sec. 5 which indicates the systemic velocity may differ from that. The contribution of the Fe II template and H I line profiles are shown in Fig. 2.
@@ 3 Emission lines: inferences and comparisons
From the spectral fit described above, we calculate a line luminosity111Adopting km s-1 Mpc-1, , and ; and using the calculation provided by Wright (2006). of erg s-1, and can extrapolate the optical continuum luminosity to be erg s-1 at 5100 Å. The equivalent width of H is Å, significantly higher than the typical range of Å found for most QSOs (Osterbrock, 1978; Hu et al., 2008).
The line ratio H/H is lower than the value of 0.47 for case B recombination at K (Osterbrock, 1989). It is also marginally lower than the range of values reported by La Mura et al. (2007) for broad line AGN, although it is consistent with the lowest values these authors found for several of their sample. We note that H is within our spectral range, and we find222the S/N of the line ratios is relatively high, in this case , because their shapes are tied to be the same. /H which is slightly less than the lowest values reported by the same authors. As such the line ratios appear physically plausible, and indicate that there might be some modest extinction towards the broad lines.
For the full line profile, we find FWHM km s-1 and km s-1 so that FWHM/. This ratio, has been discussed significantly in the literature (Collin et al., 2006) and is known to vary between AGN. The value we find is significantly lower than the canonical value of 2 often adopted for local AGN when deriving scaling relations (Collin et al., 2006; Woo et al., 2015), and is a direct result of the blue wing on the line profile.
We also calculate the ratio333An alternative that is also commonly used is the ratio of the equivalent widths (Boroson & Green, 1992; Shen & Ho, 2014). In practice it makes little difference because the continuum flux density changes very little between the corresponding wavelengths. where is the flux of Fe II in the range 4434-4684 Å (Du & Wang, 2019). As noted above, this is indicative of the accretion rate and, in a sample of AGN selected to be super-Eddington, Du & Wang (2019) find many of their objects have . Despite the high luminosity of J0529, our value of is modest compared to that sample.
3.1 Single epoch estimates
@@
3.1.1 BLR size from scaling relations

### v2

@@
We present the first near-infrared interferometric data of a QSO at z=4. The K-band observations were performed with GRAVITY+ on the VLTI using all four UTs, detecting a differential phase signal that traces the spatially resolved kinematics for both the H and H lines in the broad line region. We fit the two lines simultaneously with an updated model that includes distinct rotating and conical outflowing components. For the best fit model, more than 80% of the H I line emission from the BLR originates in an outflow with a velocity up to km s-1. This is oriented so that our line of sight is along an edge of the conical structure, which produces the prominent blue wing on the line profile. A combination of anisotropic line emission and mid-plane opacity lead to the single-sided phase signal. The model is able to qualitatively match both the outflowing C IV line profile and the systemic O I fluorescent emission. The derived black hole mass of M⊙ is the highest redshift black hole mass measurement to date obtained directly from BLR dynamics. It is an order of magnitude lower than that inferred from various single epoch scaling relations, and implies that the accretion is highly super-Eddington. With reference to recent simulations, the data suggest that this QSO is emitting close to its radiative limit in a regime where strong outflows are expected around a polar conical region.
@@
Galaxies: active – quasars: individual: SMSS J052915.80-435152.0 – quasars: emission lines – quasars: supermassive black holes – Infrared: galaxies1 Introduction
@@
The cosmic epoch at marks a key period between the end of reionisation at (rob22; loe01) and the most active phase of galaxy formation at cosmic noon around (for20; tac20). There is an increasing interest in spatially resolving the molecular and ionised emission lines in galaxies on kiloparsec scales during this epoch (fev20; wyl22; gal23; her25; ubl24), because doing so can provide insights into the processes that have shaped the universe into what it is now. The role of active galactic nuclei (AGN) is equally important because the black hole accretion rate is rising rapidly during this epoch until it peaks around the same time as galaxy formation (air15; kim24). Indeed, both the major impact that feedback from AGN can have on galaxy growth (fab12; kin15; fan23), and the existence of local relations between black hole mass and their host galaxy properties (fer00; geb00; gul09; mcc13; sag16), indicate that supermassive black holes should co-evolve with galaxies (kor13; hec14).
@@
Observations with JWST, as nicely illustrated by ada25, have recently highlighted a population of AGN with luminosities overlapping those of local Seyfert nuclei, spanning the entire redshift range from cosmic noon to reionisation (koc23; ubl23; gre24; mai24; mat24; suh25; juo25; sch25). The data have unexpectedly indicated that, compared to local relations, these black holes are overmassive with respect to their host galaxies. The implication is that they must have grown very rapidly, requiring sustained near-Eddington, or even super-Eddington, accretion, depending on the mass of the seed black holes (kok23; lar23; bog24; jeo24; mai24; meh24; suh25). However, concerns have been raised about the scale of uncertainties on both the black hole mass and host galaxy mass estimates, and on the use of scaling relations to derive black hole masses (lam24; lup24; ada25; ber25).
@@
Complementary to these efforts are studies focussing on highly luminous QSOs, which are expected to drive the powerful outflows that have the greatest impact on host galaxy evolution (bis17). Such QSOs are inferred to have black hole masses M⊙, with a significant fraction close to M⊙, and so are radiating at around their Eddington luminosity (vie18). The mass outflow rates derived from the prominent broad [O III] lines exceed 1000 M⊙ yr-1, and have kinetic energies around 1% of the bolometric luminosity. vie18 showed that they follow the expected trend (sul11; sul15), where those exhibiting prominent [O III] have moderately strong C IV emission with modest blue-shift, while QSOs without such strong [O III] emission have weaker C IV that is more strongly blue-shifted. These authors argued that this could be due to orientation effects affecting observation of the outflows in the broad line region (BLR); and comparison of the C IV blue-shift and FWHM led them to propose that the line profile is a combination of virialised and polar outflowing components. Extremely luminous QSOs at have also been targeted by mid-infrared JWST observations with the aim of resolving the host galaxy (wyl22; vay24), The data have revealed extended [O III] emission in the host galaxies, the outflows, and the circumgalactic medium.
@@
While the properties of outflows on large scales correlate with AGN luminosity, there are also significant departures, for example due to the short timescale variability of AGN which can lead to a wide range of mass loading factors (zub20). Similarly, how the outflow is driven, the relative importance of different gas phases, the role of dust, and the degree of clumpiness, can all impact the outflow properties and may change with distance from the AGN (kin11; fau12; zub12; som15; war24). As such, resolving structures in these massive outflows nearer to their launching sites is central to understanding them, but doing so requires additional techniques due to the extremely small scales involved. Velocity resolved reverberation mapping (RM) of nearby AGN has shown that there can be an outflow component to the BLR (den09; gri13; du16a). More recently, near-infrared interferometry with GRAVITY has been used to spatially resolve the BLR (gra18stu; gra20sha; gra21shi), finding that in some local AGN, outflows can dominate the BLR kinematics (gra24san). A QSO at has now been observed with the same technique, resolving the BLR kinematics in a highly accreting QSO (abu24). Observations of several more luminous QSOs at the same epoch have revealed clear signatures of radial flows in a significant fraction of their BLRs (GRAVITY+ Collaboration et al., in prep).
@@
In this paper we take the next step towards higher redshifts, and present interferometric data from GRAVITY+ that resolve the BLR in SMSS J052915.80-435152.0 (hereafter J0529), a highly luminous QSO at z=3.96 (wol24). Our focus is two-fold. One aspect is to probe the AGN driven outflow on sub-parsec scales, where a broad-line region outflow may be launched. The other is to make a dynamical estimate of the black hole mass, as part of a larger effort to understand whether the scaling relations derived locally can be reliably applied at high redshifts and if needed re-calibrate those relations. Our observations are presented in Sec. 2. We provide a context for our analysis in Sec. 3 by reporting the BLR size and BH mass expected from scaling relations. We then describe our BLR model in Sec. 4 and present our fitting results in Sec. 5, speculating about why they differ so significantly from the scaling relations, and showing how the model can be used to match other emission lines. We summarise our results in Sec. 7.
@@ 2 Observations and Data Reduction
The data analysed here were obtained with GRAVITY (grav17) as part of programme 114.27UV. They make use of the dual field mode of GRAVITY-wide (grav22) and the new adaptive optics modules (grav25), which are upgrades towards GRAVITY+. As used here, GRAVITY combines the light of all four 8-m UTs of the VLT, to probe spatial scales of milli-arcsec and perform spectro-astrometry at 10 micro-arcsec levels.
The data were obtained on five nights from Sep. to Dec. 2024 with typical seeing (at zenith, and at 500 nm) on different nights of 0.5″ and 0.7″, and coherence timescales of ms. The MEDIUM-COMBINED dual-field wide mode was used, providing medium () spectral resolution and combining the polarisations, and with more than 2″ between the science target and fringe tracking reference star. A total of almost 7.8 hr integration was obtained in 64 exposures as shown in Table 1. After processing the data with the GRAVITY instrument pipeline (v. 1.7.0b1), particularly noisy data sets were rejected, reducing the number used from 64 to 59. An additional step was also made to compensate the wavelength dependent variability caused by the dispersion of air in the non-vacuum delay lines as described in Appendix A of gra20sha, using updated components calibrated from GRAVITY-wide observations. Due to the limited signal-to-noise of individual science frames, it was applied on a night-by-night basis: the scaling derived from the combined frames on each night (excluding spectral regions near the H and H lines) was applied to each of those frames individually. Although this approximation cannot take into account changes in delay line length due to tracking, it appears to be sufficient in these circumstances. The resulting differential phase spectra from all nights were then combined with equal weighting, after rejecting (at most a few) deviant values in each spectral channel.
@@
| 27 Sep 2024 | 0.47 | 8.4 | 3 | 4300 |
| 14 Nov 2024 | 0.72 | 5.9 | 20 | 4100 |
| 15 Nov 2024 | 0.48 | 10.4 | 22 | 4100 |
| 12 Dec 2024 | 0.53 | 5.9 | 6 | 4100 |
| 14 Dec 2024 | 0.72 | 4.7 | 13 | 4100 |
An additional K-band spectrum was obtained with the ERIS adaptive optics imager and spectrometer (dav23) as part of programme 112.25M3. ERIS includes an integral field spectrometer (SPIFFIER) which was used here in seeing limited conditions (0.71″ at zenith and 500 nm) with the largest 8″ field of view, and the K-band grating covering 1.96-2.45 m at a resolution of . The data were taken as 11 dithered exposures of 300 s, yielding a total integration time of 55 mins, and reduced with the ERIS instrument pipeline (v 1.6.4). Verification of the wavelength calibration was performed using OH sky emission lines. Telluric correction and flux calibration were performed using the B9 IV standard star HD 38343 which has K=7.87 mag. Following this, additional steps were performed to interpolate over the noisiest pixels in the datacube and reduce wavelength dependent variations in the background level. The flux density of the QSO was measured from a spectrum extracted in a large 1.25″ aperture, yielding K = 13.88 mag.
@@ 2.1 Image
An image was extracted from the datacube by summing channels in the range 2.028–2.290 m, indicating that the spatial resolution achieved on the QSO itself in the K-band was 0.43″. The image is shown in Fig. 1 on a logarithmic scaling to highlight faint sources. The QSO shows no sub-structure, and the only other source in the field is the M dwarf star previously noted by wol24 a few arcsec north of the QSO. This strengthens the assessment made by those authors that there is no sign of strong lensing, and hence the luminosity is intrinsic.
@@ 2.2 Spectral decomposition
A final spectrum with higher signal-to-noise (S/N) of 85 (but slightly lower flux) than that above was extracted in a 0.62″ aperture better matched to the FWHM. This is used to recover the spectral profile of the broad H I lines, which requires decomposition of the spectrum in order to deblend them from the Fe II features which are known to be strong in highly accreting QSOs (bor92; du16b). Since it is already clear from Fig. 2 that the H I lines have a prominent blue wing, we fit these with a pair of Gaussians. This is done solely to match the overall profile and we do not consider the Gaussians individually. We use the same velocity profile for both H I lines, which, although not physically required, is necessary due to limitations of the data. In addition we required zero velocity offset between the H I lines, which was an important constraint on the fit. The reason is that the blue wing of the H line is strongly blended with Fe II and so is primarily constrained by the H line. On the other hand the red side of the H line does provide a primary constraint, whereas the H line is affected by both atmospheric absorption (leading to narrow features and also affecting the slope) and potentially some emission from a blue-shifted wing of [O III] line emission, the peak of which would be just out the spectral range. While we do not have direct evidence for [O III] (and indeed the [O III] line at 4363 Å is not detected although it is expected to be a factor 10-100 weaker), in luminous high-z samples it can be associated with an outflow as indicated by the strong blue-shifted C IV emission described in Sec. 6 (mar17). The intrinsic continuum is expected to be a power law, but over this narrow wavelength range we have adopted a linear function. We consider several templates for the optical Fe II emission, including those based on I Zw 1 (ver04), Mrk 493 (par22), a composite based on several QSO spectra (tsu06), and an adaptable template derived from multiple AGN (kov10). After trying these options, we adopted the template from par22 because it provides the best match to the Fe II emission on the blue side of the H line; in other regions the templates were equivalently good. Although the properties of the Fe II complex were not tied to the Balmer lines, the broadening of km s-1 FWHM is comparable to the width of the Balmer lines as one would expect if they arise in similar locations (hu15). Finally, specific features in the spectrum that were clearly due to atmospheric effects were excluded during the fitting process. The resulting fit has . The observed spectrum and its model fit, showing the contribution of the Fe II template and H I line profiles, are shown in Fig. 2.
@@ 3 Emission lines: inferences and comparisons
From the spectral fit described above, we calculate a line luminosity111Adopting km s-1 Mpc-1, , and ; and using the calculation provided by wri06. of erg s-1, and can extrapolate the optical continuum luminosity to be erg s-1 at 5100 Å. The equivalent width of H is Å in the observed frame and hence Å in the rest frame, putting it within the typical range of Å found for most QSOs (ost78; hu08).
The line ratio H/H is lower than the value of 0.47 for case B recombination at K (ost89). It is also marginally lower than the range of values reported by mur07 for broad line AGN, although it is consistent with the lowest values these authors found for several of their sample. We note that H is within our spectral range, and we find222the S/N of the line ratios is relatively high, in this case , because their shapes are tied to be the same. /H which is slightly less than the lowest values reported by the same authors. As such the line ratios appear physically plausible, and indicate that there might be some modest extinction towards the broad lines.
The full line profile has FWHM km s-1 and km s-1 so that FWHM/. This ratio, has been discussed significantly in the literature (col06) and is known to vary between AGN. The value we find is significantly lower than the canonical value of 2 often adopted for local AGN when deriving scaling relations (col06; woo15), and is a direct result of the blue wing on the line profile.
We also calculate the ratio333An alternative that is also commonly used is the ratio of the equivalent widths (bor92; she14). In practice it makes little difference because the continuum flux density changes very little between the corresponding wavelengths. where is the flux of Fe II in the range 4434-4684 Å (du19). As noted above, this is indicative of the accretion rate and, in a sample of AGN selected to be super-Eddington, du19 find many of their objects have . Despite the high luminosity of J0529, our value of is modest compared to that sample.
@@
3.1 BLR size from scaling relations